OpenStudy (anonymous):

1)square rt x-10+4=0 answer is 36 but I need help explaining how I got it. I'm helping my 12 year old son and I'm old fashion calculator is my thing....please help 2) 3x4/5-6=9

OpenStudy (anonymous):

your question is not clear. repost it, using brackets to clarify what you mean. does the square root apply to everything? or just the first term? or just the first two terms?

OpenStudy (anonymous):

\[\sqrt{?}\] x-10+4=0 3x 4/5-6=9

OpenStudy (anonymous):

again, not clear. rewrite it properly. or at least answer my previous question

OpenStudy (anonymous):

or there is an option to draw. use the draw button below to draw your question. | | v

OpenStudy (anonymous):